What is the period of the wave motion for a wave with a frequency of 0.36 kHz
1
Expert's answer
2011-10-27T09:46:12-0400
Period T=1/f where f - frequency
o.36 kHz=0.36*10^3 Hz
T=1/(0.36*10^3)=0.0028 c